OpenStudy (anonymous):

WILL SOMEONE PLZZ HELP ME WITH A FEW MATH TESTS? ITS FOR SCHOOL AND I JUST CANT DO MATH FOR THE LIFE OF ME BUT I NEED TO DO THESE TO BUMP MY GRADE TO PASS. PLZZZ HELP!!!!!!!!!!!!!

OpenStudy (anonymous):

please less caps =) and yes lol

OpenStudy (anonymous):

thanks so much lol sorry

OpenStudy (anonymous):

it ok =)

OpenStudy (anonymous):

so what is the question ...

OpenStudy (anonymous):

What is the solution to the system of equations? 4x+y=2 x-y=3 (_____, _____) ?

OpenStudy (anonymous):

@jakashaka123

OpenStudy (anonymous):

nvm I think I might have got the first test but can you help me with some more in a few minutes plz?

OpenStudy (anonymous):

yeah i can i was doing some stuff srry

OpenStudy (anonymous):

Its okay thanks ill brb

OpenStudy (anonymous):

aight =)

OpenStudy (anonymous):

Solve using the substitution method 4x + 2y=11 x-2= -2y A: (3, -1/2) B: (0, -2) C: (0, -11/2 D: (7, -5/2) @jakashaka123

OpenStudy (anonymous):

@rational

OpenStudy (allieeslabae):

I can help just give me a moment!

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

well your x's are on left and y's are on right

OpenStudy (anonymous):

Wouldn't it be A?

OpenStudy (anonymous):

for a you have 4*3+-1/2(2) (hint-which is one =))

OpenStudy (anonymous):

so a is your answer =)

OpenStudy (anonymous):

yes =)

OpenStudy (anonymous):

its substitution not make my life a living nightmare =)

OpenStudy (anonymous):

Okay can you guys help me with another one?

OpenStudy (anonymous):

yesh =) i might be able to =)

OpenStudy (anonymous):

i have 2 minuets

OpenStudy (anonymous):

1 now =)

OpenStudy (anonymous):

nm i gtg bye guys / girls

OpenStudy (anonymous):

bye @Allieeslabae

OpenStudy (anonymous):

add me on kik i can answer on there too so =)

OpenStudy (anonymous):

its insanegamer23

OpenStudy (allieeslabae):

Hey its B! okay bye!

OpenStudy (anonymous):

Solve using the substitution method y+2x=7 14-4x=2y A: The solution is (1,5) B: The solution is (21,0) C.There are an infinite number of solutions. D; There is no solution.

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

ok so lets start with 1+2*5=

OpenStudy (anonymous):

thats to see if a is correct

OpenStudy (anonymous):

its 11

OpenStudy (anonymous):

but we need 7 soo

OpenStudy (anonymous):

21+0=7

OpenStudy (anonymous):

so its not b either =)

OpenStudy (anonymous):

no it is a nm

OpenStudy (anonymous):

5 is y and 1 is x